### v1

# Prompt to Restore, Restore to Prompt: Cyclic Prompting for Universal Adverse Weather Removal

Abstract
Universal adverse weather removal (UAWR) seeks to address various weather degradations within a unified framework. Recent methods are inspired by prompt learning using pre-trained vision-language models (e.g., CLIP), leveraging degradation-aware prompts to facilitate weather-free image restoration, yielding significant improvements. In this work, we propose CyclicPrompt, an innovative cyclic prompt approach designed to enhance the effectiveness, adaptability, and generalizability of UAWR. CyclicPrompt Comprises two key components: 1) a composite context prompt that integrates weather-related information and context-aware representations into the network to guide restoration. This prompt differs from previous methods by marrying learnable input-conditional vectors with weather-specific knowledge, thereby improving adaptability across various degradations. 2) The erase-and-paste mechanism, after the initial guided restoration, substitutes weather-specific knowledge with constrained restoration priors, inducing high-quality weather-free concepts into the composite prompt to further fine-tune the restoration process. Therefore, we can form a cyclic “Prompt-Restore-Prompt” pipeline that adeptly harnesses weather-specific knowledge, textual contexts, and reliable textures. Extensive experiments on synthetic and real-world datasets validate the superior performance of CyclicPrompt. The code is available at: https://github.com/RongxinL/CyclicPrompt.
Index Terms:
Universal adverse weather removal, Image restoration, Cyclic prompt
I Introduction
Images captured in adverse weather, such as rain, snow, and fog/haze, often suffer from diminished quality, which presents significant challenges for outdoor vision systems [1, 2, 3]. Most methods primarily focus on crafting specialized networks tailored to certain weather conditions, including image deraining [4, 5, 6], defogging/dehazing [7, 8], and desnowing [9]. Some other studies [10, 11, 12] have developed a general restoration architecture and train discrete copies independently to accommodate different weather conditions. Although notable outcomes have been obtained in their intended scenarios, these task-specific models struggle to generalize effectively across different weather types and even intensity levels. This implies that tedious specialized weights are necessitated in practical applications, increasing computational costs and resources. Chen et al. [13] develop pre-trained image processing models comprising multiple heads aligned to different low-level tasks, which excavates the model capability to deal with different corruptions but requires further fine-tuning on each specific dataset.
Recent studies [14, 15, 16, 17, 18, 19] have shifted towards developing generic solutions to address multiple adverse weather challenges within a single, unified model. Although some generalizability improved on synthetic datasets, their adaptability to real-world scenarios remains limited. Inspired by the power of large-scale pre-trained vision-language models (PVLM) [20] in various computer vision tasks [21, 22], some works leverage off-the-shelf PVLMs to extract degradation-aware prompts in linguistic [23, 24] or visual [25, 26] domains and interact with the main feature flow hierarchically. For instance, DA-CLIP [23] aligns text embeddings with visual representations to predict the degradation prompt integrated into other restoration models. Liang et al. [24] construct a question prompt based on pre-trained LISA [21] to estimate the weather types and levels. Despite remarkable advancements, existing prompt-driven methods for UAWR still fall short in several limitations: 1) While PVLMs play a pivotal role in navigating weather-related content, describing weather types through simple text or converted image embeddings remains challenging, particularly in scenarios with complex weather (e.g., rain+fog, Fig. 1). 2) Visual prompts from corrupted images may not adequately express the spatial context or even introduce ambiguous semantics due to the interference of weather degradations.
Figure 1: Our CyclicPrompt is capable of removing multiple adverse weather degradations in one model and producing more visually pleasant images than several state-of-the-art UAWR methods: TransWeather[15], WeatherDiff[27], and T3-Diffweather[26]. The quantitative comparisons in PSNR and SSIM also indicate the preferable performance of our method obviously.
To tackle these issues, this paper presents an in-depth investigation into the construction of effective prompts to fully unleash the potential of prompt learning. To this end, we propose a remarkably different cyclic prompt learning approach for UAWR, termed CyclicPrompt, which integrates weather-specific knowledge, textual context, and reliable textures, so that better results can be yielded. Specifically, we construct a composite context prompt (C2P) to perceive weather-related information and context-aware representations. The first component of C2P aligns with existing methods that extract CLIP-based visual features but is further enhanced by incorporating learnable input-conditional vectors, enabling dynamic adaptation of weather-specific knowledge to improve adaptability. The second is derived from explicit textual descriptions to mitigate the semantic ambiguities caused by adverse weather. While the composite prompt shows evident effects, it may not sufficiently facilitate more realistic texture recovery as it primarily focuses on compact semantics rather than visual characteristics. To overcome this, after the C2P-driven UAWR, we propose the erase-and-paste mechanism (EPM). It leverages constrained restoration priors to substitute the weather-related part with high-quality weather-free concepts in C2P, strengthening the information flow in a cyclic “Prompt-Restore-Prompt” process. On the other hand, along with EPM, we also devise the residual prior modulator based on restored results to provide plausible texture clues and structural appearance with different granularities.
In summary, the main contributions of our work are
  • We present CyclicPrompt, a distinct cyclic prompt-driven universal framework that effectively and adaptively removes weather degradations in an all-in-one manner for UAWR. Extensive experiments on both synthetic and real-world datasets demonstrate its superiority.
  • We construct a composite context prompt (C2P) that combines CLIP-equipped weather-specific knowledge, learnable input-conditional vectors, and textual prompt to capture weather-related information and context-aware representations.
  • We introduce the erase-and-paste mechanism (EPM), leveraging constrained restoration priors to embed weather-free representations into the C2P, thereby refining the restoration process. Additionally, the residue prior modulator is designed to enrich visual characteristics for better detail recovery along with EPM.
The rest of the paper is organized as follows. Section II reviews the related works. Section III introduces the proposed CyclicPrompt in detail. Section IV shows the experimental results of our method and validates the effects of the proposed components. In Section V, we conclude this work.
II Related Work
II-A Task-Specific Adverse Weather Removal
Task-specific methods [4, 8, 10, 11] train the individual model for each specific weather type like rain [4, 5, 28, 29], snow [30, 9, 31], and fog [8, 32, 33, 34]. In literature, tremendous methods are tailored to these degradations, known as image draining, dehazing, and desnowing et al.. Here, we provide a brief review of them.
Image Deraining. Rainy weather usually affects the images in two ways: rain streaks and raindrops. Fu et al. [35] propose the first convolutional neural network (CNN)-based method DerainNet for rain streak removal. As motivated, serial methods are developed that utilize advanced deep networks [36, 37, 38, 39, 40, 41, 42] or rain-related physical modeling [33, 43, 44, 45] to pursue promising results. To alleviate the limitation of CNNs in long-range dependency capturing, [5, 46, 47] combines the locality of CNNs and the global capacity of transformers, which performs favorably. Raindrops adhere to glass surfaces (e.g. windows or camera lens) and occlude the regions of background scenes. To mitigate the damage of visibility, a popular approach [48, 29, 49] is to detect the raindrop masks from rainy inputs to help recover raindrop-free ones. There are also several methods [28, 50] jointly remove the rain streaks and raindrops, significantly improving the applicability in real rainy scenes.
Image Dehazing. Hazy images often suffer from color distortion, blurry textures, and poor quality. In past years, the prevailing progress of image dehazing has been witnessed in deep learning [51, 52, 53, 54, 55, 56]. DehazeNet [51] introduces the pioneering CNN for haze removal. Considering that the haze on objects is highly related to depth, recent research [57, 58, 59] recover from the hazy image by depth estimation and depth-aware reconstruction. Tian et al. [60] integrate diffused texture priors for high-quality image restoration, which has demonstrated powerful generative ability.
Image Desnowing. Unlike rain or haze, the particle shapes and sizes in snow introduce greater complexity. From the first deep learning method DesnowNet [9], deep learning has dominated this task [30, 61, 62, 63]. DDMSNet [31] learns semantic- and geometry-aware representations utilizing semantic and depth priors. In [64], Chen et al. propose the CPLFormer that learns from cross-scale snow prototypes to handle complicated snow landscapes. Zhang et al. [65] propose the hybrid color space-based desnowing network based on the analysis of snowflake noises in the Hue channel. While the above studies acquire impressive achievements on specific weather conditions, they ignore other adverse weather which limits their scope of practical applications.
II-B Universal Adverse Weather Removal
UAWR [15, 66] aims to develop a single model to recover weather-free images under different adverse weather, which is a long-standing problem in the computer vision community. Many of them aggregate parallel weather-specific encoders [14, 18, 17] or decoders [67] where each learns certain weather representations in a unified architecture for multi-weather removal. AWRCP [19] utilizes the codebook priors by pre-trained VQGAN [68] from undistorted images to improve recovered details. Motivated by the generative capacity of diffusion models [69, 22], WeatherDiff [27] presents a patch-based diffusive algorithm, achieving promising generalization ability on synthetic and real-scenario multi-weather restoration. In more recent years, prompt learning with pre-trained models [20, 21] has emerged as a powerful tool to learn degradation-aware information effectively from degraded images [23, 25, 24], thereby empowering the effectiveness and generalizability of restoration models. This work introduces a cyclic prompt learning framework for UAWR, which shows the superiority of various weather degradations.
Figure 2: Overview of the proposed CyclicPrompt built upon a transformer architecture following a cyclic “Prompt-Restore-Prompt” process for UAWR. We first construct a composite context prompt (C2P) containing weather-specific knowledge, input-conditional vector, and textual prompt to perceive weather-related information and context-aware representations. After the C2P-driven UAWR, the erase-and-paste mechanism (EPM) leverages constrained restoration priors to substitute the weather-related part in C2P with the weather-free prompt to facilitate the decoder flow further. The residual prior is extracted to provide texture clues and structural appearance for better detail recovery.
II-C Prompt Learning
Prompt learning originates from natural language processing (NLP) [70, 71], which provides in-context information to adapt models to other tasks with minimal fine-tuning [72, 73], particularly benefiting large-scale pre-trained models. Inspired by the pioneering attempt CLIP [20] that applies PVLMs to downstream tasks, prompt learning has been expanded in computer vision domains [74, 75, 76], which incorporates handcrafted prompt templates or learnable prompt tokens to guide the interaction between textual and visual tokens. CoOp [77] models a prompt’s context words with learnable vectors to adapt CLIP-like vision-language models for downstream image recognition. CoCoOp [78] further introduces conditional context optimization that learns input-conditional dynamic prompts with the static prompts in CoOp [77] to improve its transferability. Building upon these advancements, researchers have also explored the advantages of prompt learning for low-level vision tasks. CLIP-LIT [79] devises iterative prompt learning to characterize backlit and well-lit images more precisely. SeeSR [80] generates image tag text as a semantic prompt for real-world image super-resolution. This work exploits prompt learning but induces significant insights to achieve effective and generalized UAWR.
III Method
III-A Overview
The overall framework of the proposed CyclicPrompt is illustrated in Fig. 2. The core components are 1) the composite context prompt (C2P) capturing weather-related information and context-aware textual representations (Sec. III-B), and 2) the erase-and-paste mechanism (EPM) providing weather-free representations into the C2P for further refinement (Sec. III-C). We build the prompt block to adapt the deep image features based on our C2P (Sec. III-D). Moreover, we also extract the residual prior to improve the detail recovery (Sec. III-D). Our CyclicPrompt establishes a “Prompt-Restore-Prompt” pipeline with the C2P and EPM.
Given a low-quality image $\mathbf{I}_{LQ}$ degraded by unknown adverse weather, the goal of our CyclicPrompt is to learn a universal model $\mathcal{M}$ and restore a photo-realistic weather-free image. We build the overall framework upon a U-shaped transformer architecture. Generally, the transformer encoder first maps the high-resolution input $\mathbf{I}_{LQ}$ into a lower dimensional latent space via $\mathbf{F}_{e}=\mathcal{E}(\mathbf{I}_{LQ})$. The transformer decoder gradually recover the clean output $\mathbf{I}_{HQ}$ from $\mathbf{F}_{e}$. We incorporate C2P $\mathbf{P}_{c}=\{\mathbf{P}_{v},\mathbf{P}_{t}\}$ in the decoder part via the prompt block (Fig. 3) to guide the decoding process. After the initial restoration, the EPM replaces the weather-related information $\mathbf{P}_{v}$ with the weather-free prompt $\mathbf{P}_{w}$ extracted from $\tilde{\mathbf{I}}_{HQ}$, constituting an enhanced prompt $\hat{\mathbf{P}}_{c}$ to strengthen the information flow. Simultaneously, we also extract the residual prior $\mathbf{R}$ from $\tilde{\mathbf{I}}_{HQ}$, which is then equipped with a residual prior modulator (RPM) in the prompt block to provide detail representations for restoration fidelity. We take $\mathbf{R}$ and $\hat{\mathbf{P}}_{c}$ as the guidance to facilitate the decoding process further, endowing the model with more promising generative capacity and thus producing the final reconstruction $\hat{\mathbf{I}}_{HQ}$.
In CyclicPrompt, both two restoration processes are constrained by L1 loss. Therefore, the total loss function for training our model can be formulated as
$\mathcal{L}=\|\tilde{\mathbf{I}}_{HQ}-\mathbf{I}_{HQ}\|_{1}+\|\hat{\mathbf{I}}_{HQ}-\mathbf{I}_{HQ}\|_{1},$ (1)
where $I_{HQ}$ denotes the ground-truth image of $\mathbf{I}_{LQ}$.
III-B Composite Context Prompt
In this work, we consider the prompt in the UAWR task from two aspects to build the C2P $\mathbf{P}_{v}$. First, we integrate weather-specific knowledge $\mathbf{P}_{k}$ and learnable input-conditional vector $\mathbf{P}_{i}$ as the visual prompt $\mathbf{P}_{c}$ to perceive the weather-dedicated information. Second, the text description for the input image can be easily accessible to encapsulate the semantic information, serving as the textual prompt $\mathbf{P}_{t}$.
Recent works [23, 80] have demonstrated the potential of text representations in image restoration as it is easily accessible and intuitively expresses the semantics of the objects, environment, and contexts in an image. In this work, for the input image $\mathbf{I}_{LQ}$, we apply BLIP [81] to generate corresponding caption $\mathbf{T}$. Then, considering CLIP [20] can effectively encode the correlations between visual and linguistic data, we obtain the textual prompt $\mathbf{P}_{t}$ by a frozen CLIP text encoder as follows:
$\mathbf{P}_{t}=\mathcal{E}_{clip}^{text}(\mathbf{T}),$ (2)
where $\mathcal{E}_{clip}^{text}$ denote the text encoding operation. $\mathbf{P}_{t}\in\mathbb{R}^{1\times D}$ and $D$ is the channel dimension.
In real scenarios, where complex degradation issues are often encountered, accurately modeling the degradation through text descriptions remains a significant challenge. To alleviate this limitation, we exploit the frozen CLIP image encoder to transform $\mathbf{I}_{LQ}$ into deep features and perceive weather degradations, which are then processed by a multilayer perceptron (MLP), yielding the weather-specific knowledge $\mathbf{P}_{k}\in\mathbb{R}^{1\times D}$
$\mathbf{P}_{k}={\rm{MLP}}_{1}(\mathcal{E}_{clip}^{img}(I_{LQ})),$ (3)
where $\mathcal{E}_{clip}^{img}$ denotes the image transformation function. To improve the adaptability of models, inspired by the conditional prompt learning in CoOp [77], we incorporate a learnable input-conditional vector $\mathbf{P}_{i}$ for the CLIP-equipped image feature to dynamic calibrate and adapt the weather-dedicated information to complex unknown weather conditions, which is combined with $\mathbf{P}_{k}$ to generate the visual prompt
$\mathbf{P}_{v}=\mathbf{P}_{k}+\mathbf{P}_{i},$ (4)
where $\mathbf{P}_{i}\in\mathbb{R}^{N\times D}$ is random initializing tunable vector which length is $N$. We replicate $\mathbf{P}_{k}$ to $N$ copies, thus being the same size with $\mathbf{P}_{v}$.
Based the above discussion, the initial C2P $\mathbf{P}_{c}\in\mathbb{R}^{(1+N)\times D}$ can be constructed by concatenating the weather-related visual prompt $\mathbf{P}_{v}$ and the textual prompt $\mathbf{P}_{t}$
$\mathtt{Initial:}\quad\mathbf{P}_{c}={\rm{Concat}}(\mathbf{P}_{v},\mathbf{P}_{t}).$ (5)
We inject C2P into the CyclicPrompt framework at the transformer decoder through prompt blocks (Fig. 3) to interact with the mainstream image features to improve the representations for the first guided restoration process (1st iteration) in our cyclic prompt learning framework, i.e. “Prompt-Restore”.
III-C Erase-and-Paste Mechanism
With the guidance of C2P, we have the initial result $\tilde{\mathbf{I}}_{HQ}$. In contrast to existing methods [82, 24], this work makes a forward step, which regards the recovered image $\tilde{\mathbf{I}}_{HQ}$ as a restoration prior to distill restoration-related information for further refinement. On the one hand, $\tilde{\mathbf{I}}_{HQ}$ is fed into another frozen CLIP image encoder followed by a trainable MLP to extract high-quality image features that are referred to the weather-free prompt $\mathbf{P}_{w}\in\mathbb{R}^{1\times D}$
$\mathbf{P}_{w}={\rm{MLP}}_{2}(\mathcal{E}_{clip}^{img}(\tilde{\mathbf{I}}_{HQ})).$ (6)
We propose the EPM, which first erases the weather-related representations $\mathbf{P}_{v}$ in $\mathbf{P}_{c}$ and then pastes $\mathbf{P}_{w}$ on it,
$\mathbf{P}_{v}\rightleftarrows\mathbf{P}_{w},$ (7)
$\mathtt{Cyclic:}$ $\hat{\mathbf{P}}_{c}={\rm{Concat}}(\mathbf{P}_{w},\mathbf{P}_{t}),$
where $\hat{\mathbf{P}}_{c}\in\mathbb{R}^{2\times D}$ denotes the new reconstituted cyclic C2P. Therefore, there is a 2nd iteration, “Restore-Prompt”, which applies $\hat{\mathbf{P}}_{c}$ as the guidance that induces weather-free concepts to optimize the restoration features, empowering the model with more effectiveness and generalizability.
Besides, in EPM, we implement residual prior extraction on the initial restored image $\tilde{\mathbf{I}}_{HQ}$ by subtracting its minimum channel from the maximum channel, i.e. residue channel, which has shown effects on rain removal [14, 42]. Such a prior $\mathbf{R}$ is weather-invariant and allows maximized maintaining the contexts. This process can be formulated by
$\mathbf{R}=\max_{c\in\{r,g,b\}}\tilde{\mathbf{I}}_{HQ}-\min_{c\in\{r,g,b\}}\tilde{\mathbf{I}}_{HQ},$ (8)
where $\max_{c\in\{r,g,b\}}$ and $\min_{c\in\{r,g,b\}}$ denotes the maximum and minimum RGB channel of $\tilde{\mathbf{I}}_{HQ}$. To take full advantage of the residual prior, in the 2nd cyclic prompt-based restoration, it complements with $\hat{\mathbf{P}}_{c}$ in the prompt block (see Fig. 3) to enrich the texture clues and structural appearance.
By combining C2P and EPM, we can form a remarkedly distinctive cyclic prompt learning-based framework, providing holistic representations such that the restoration process can be effectively adapted for adverse weather.
Figure 3: The 4-level transformer architecture in CyclicPrompt and each level contains several transformer blocks. The prompt block is inserted before each level of the decoder, where each contains two branches with the branch selection controlled by the restoration process: the 1st iteration with initial C2P (top branch) and the 2nd with cyclic C2P (bottom branch). We use the cross-hierarchy information mining blocks (CHIMB) [83] to extract residual features.
III-D Prompt Block
We use the transformer architecture in Restormer [10] as the backbone, which adopts a 4-level transformer encoder and decoder, where each level contains several transformer blocks, illustrated in Fig. 3. The transformer encoder gradually downsamples the input image $\mathbf{I}_{LQ}\in\mathbb{R}^{H\times W\times C}$ while increasing the channel number, resulting in deep image features $\mathbf{F}_{e}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 8C}$. The prompt block is inserted before each level of the decoder. As shown in Fig. 3(a), each block consists of two branches, with the branch selection controlled by the restoration process: the 1st iteration with initial C2P (top branch) and the 2nd with cyclic C2P (bottom branch).
Specifically, let $\mathbf{X}$ denote the incoming feature fed into a certain prompt block. For the 1st iteration, the C2P $\mathbf{P}_{c}$ is projected into key and value spaces via two respectively linear layers $\mathbf{W}_{k}$ and $\mathbf{W}_{v}$, getting $\mathbf{K}\in\mathbb{R}^{N\times C}=\mathbf{W}_{k}\mathbf{P}_{c}$ and $\mathbf{V}\in\mathbb{R}^{N\times C}=\mathbf{W}_{v}\mathbf{P}_{c}$, where $\mathbf{X}$ is reshaped to the same size of $N\times C$ and then projected to the query space $\mathbf{V}\in\mathbb{R}^{N\times C}=\mathbf{W}_{q}\mathbf{X}$.We interact $\mathbf{P}_{c}$ with $\mathbf{X}$ through a cross-attention mechanism [22]
$\hat{\mathbf{X}}=\mathcal{A}((\mathbf{Q},\mathbf{\mathbf{K}}),\mathbf{V})={\rm{Softmax}}(\frac{\mathbf{Q}\mathbf{K}^{\mathrm{T}}}{d})\cdot\mathbf{V},$ (9)
where $\mathcal{A}$ denotes the cross attention operation and $\hat{\mathbf{X}}\in\mathbb{R}^{N\times C}$ is the resulted output which is then transformed to $H\times W\times C$.
When the current iteration is the second, the cyclic prompt $\hat{\mathbf{P}}_{c}$ by EPM and the residual prior $\mathbf{R}$ pass through the bottom branch sequentially. In addition to the cross attention between $\mathbf{X}$ and $\hat{\mathbf{P}}_{c}$ as the 1st iteration, there is a residual prior modulator (RPM) in the prompt block that conducts conditional modulation based on $\mathbf{R}$. As shown in Fig. 3(b), two $3\times 3$ convolutions followed by two cross-hierarchy information mining block (CHIMB) [83] are leveraged to learn informative residual feature $\mathbf{F}_{r}$ from $\mathbf{R}$ with large kernel attention and channel attention. Next, another two $3\times 3$ convolutional layers with two respective $1\times 1$ convolution layers are applied to generate a pair of affine transformation parameters ($\alpha$, $\beta$). Therefore, the residual prior can modulate the attentive features $\tilde{\mathbf{X}}$ as follows:
$\hat{\mathbf{X}}=\alpha\odot\tilde{\mathbf{X}}+\beta,$ (10)
where $\hat{\mathbf{X}}$ denotes the modulated feature and $\odot$ denotes the element-wise multiplication.
Therefore, the overall cyclic prompt learning of both iterations can be defined as
$\hat{\mathbf{X}}=\begin{cases}\mathcal{A}(\mathbf{P}_{c},\mathbf{X}),&iter.=1,\\ \mathcal{R}(\mathbf{R},\mathcal{A}(\hat{\mathbf{P}}_{c},\mathbf{X}),&iter.=2,\end{cases}$ (11)
where $\mathcal{R}$ represents the composite function of RPM. As the prompt blocks are deployed from low to high levels, one can dynamically integrate the prompt components into multi-scale features to provide sufficient guidance.
IV Experiments
We conducted extensive experiments to evaluate the effectiveness of the proposed method. This section begins by outlining the experimental settings, followed by a detailed presentation of the qualitative and quantitative results. Lastly, we perform ablation studies to analyze the contributions of each key component.
TABLE I: Quantitative Comparisons in Terms of PSNR and SSIM with State-of-the-Art Adverse Weather Removal Methods. * Denotes the Model Retrained on the All-Weather Dataset. Text Denotes the Best Performance and Underline Denotes the Second Best.
Types Methods Outdoor-Rain [33] RainDrop [29] Snow100K-S [9] Snow100K-L [9] Average
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
Rain+Fog Pix2Pix[CVPR’17] [34] 19.09 0.7100 - - - - - - - -
HRGAN[CVPR’19] [33] 21.56 0.8550 - - - - - - - -
PCNet[TIP’21] [84] 26.19 0.9015 - - - - - - - -
MPRNet[CVPR’21] [85] 30.44 0.9235 19.92 0.8385 22.13 0.7824 20.25 0.6628 23.19 0.8018
Raindrop AttentiveGAN[CVPR’18] [29] - - 31.59 0.9170 - - - - - -
RaindropAttn[ICCV’19] [48] - - 31.44 0.9263 - - - - - -
CCN[CVPR’21] [28] - - 31.34 0.9286 - - - - - -
IDT[PAMI’22] [46] 14.79 0.5416 31.87 0.9313 25.36 0.8560 21.52 0.7565 24.01 0.7763
Snow DesnowNet[TIP’18] [9] - - - - 32.33 0.9500 27.17 0.8983 - -
JSTASR[ECCV’20] [61] - - - - 31.40 0.9012 25.32 0.8076 - -
DDMSNet[TIP’21] [31] - - - - 34.34 0.9445 28.85 0.8772 - -
MPRNet[CVPR’21] [85] - - - - 34.97 0.9457 29.76 0.8949 - -
General MPRNet*[CVPR’21] [85] 29.41 0.8990 30.76 0.9100 - - 29.25 0.8803 - -
Restormer*[CVPR’22] [10] 28.54 0.8856 30.61 0.9027 - - 29.09 0.8725 - -
Uformer*[CVPR’22] [11] 26.72 0.8453 28.69 0.8819 - - 27.92 0.8476 - -
All-in-One All-in-one[CVPR’20] [14] 24.71 0.8980 31.12 0.9268 28.33 0.8820 28.05 0.9023 - -
TransWeather[CVPR’22] [15] 28.83 0.9000 30.17 0.9157 32.51 0.9341 29.31 0.8879 30.21 0.9094
TKL&MR[CVPR’22] [17] 29.92 0.9167 30.99 0.9274 34.80 0.9483 30.24 0.9020 31.49 0.9236
WeatherDiff64[PAMI’23] [27] 29.64 0.9312 30.71 0.9312 35.83 0.9566 30.09 0.9041 31.57 0.9308
WeatherDiff128[PAMI’23] [27] 29.72 0.9216 29.66 0.9225 35.02 0.9516 29.58 0.8941 31.00 0.9225
AWRCP[ICCV’23] [19] 31.39 0.9329 31.93 0.9314 36.92 0.9652 31.92 0.9341 33.04 0.9409
Tan et al.[TIP’24] [25] 30.00 0.9152 31.72 0.9325 29.85 0.8928 - - - -
Art[MM’24]  [86] 29.81 0.9088 31.54 0.9338 - - 30.61 0.9083 - -
T3-DiffWeather[ECCV’24] [26] 32.52 0.9339 32.70 0.9414 37.55 0.9641 31.11 0.9180 33.45 0.9391
GridFormer[IJCV’24] [87] 31.87 0.9335 32.39 0.9362 37.46 0.9640 31.71 0.9231 33.36 0.9392
CyclicPrompt (ours) 32.81 0.9371 32.57 0.9454 37.50 0.9655 32.16 0.9265 33.76 0.9436
IV-A Implementation Details
Following the transformer setting in [10], in CyclicPrompt, from level 1 to level 4 of the encoder, the number of transformer blocks [$B_{1}$, $B_{2}$, $B_{3}$, $B_{4}$] is set to [4, 6, 6, 8] and the channel numbers are [48, 96, 192, 384]. Since the decoder employs a two-iteration design, the number of transformer blocks is reduced by half with the same channel numbers as the encoder for computational efficiency. The prompt block is incorporated before each decoder level, totaling 3 blocks in the overall CyclicPrompt network.
We train our model with AdamW optimizer ($\beta_{1}=0.9,\beta_{2}=0.9999$, weight decay=1e-4 ) with a batch size of 4 for 800K iterations. The initial learning rate is 2e-4 and gradually reduces to 1e-6 by cosine annealing strategy. The input images are randomly cropped into $256\times 256$ during training. Randomly vertical and horizontal flips are used for data augmentation. All the models are trained on NVIDIA RTX 4090 GPUs.
Figure 4: Visual Comparisons with different UAWR methods, i.e. TransWeather [15], GrideFormer [87], T3-DiffWeather [26], WeatherDiff [27] and our CyclicPrompt on the RainDrop [29], Outdoor-Rain [33] and Snow100K-L [9] datasets.
IV-B Dataset
We conduct experiments on the All-Weather dataset [14] mainly degraded by four typical adverse weather conditions: Rain+Fog, Raindrop, and Snow. This dataset includes 18,069 images as the training data comprising 9,001 synthesized snowy images from Snow100K [9], 818 images with raindrops on windows or lenses from RainDrop [29], and 8,250 heavy rain images from Outdoor-Rain [33] for Rain+Fog. Notably, the Outdoor-Rain subset captures both rain streaks and fog-like degradation caused by rain accumulation. Due to the imbalance in the number of training samples for different weather conditions, particularly the limited size of the RainDrop subset, we perform a 10-fold resampling of the Raindrop images to ensure balanced training.
For testing, there are also 750 images from Outdoor-Rain Test1 [33] (Rain+Fog), 58 images from RainDrop TestA [29], 16,611 images from Snow100K-S [9], and 16,801 images from Snow100K-L [9]. We utilize peak signal-to-noise ratio (PSNR) and structural similarity (SSIM) as the metrics to evaluate the quantitative performance of our method.
IV-C Comparison with State-of-the-art Methods
We compare the proposed CyclicPrompt with recent state-of-the-art (SotA) adverse weather removal methods, which can be categorized into 3 groups:
  • Task-specific methods: Pix2Pix [34], HRGAN [33], PCNet [84], and MPRNet [85] for rain+fog; AttentiveGAN [29], RaindropAttn [48], CCN [28], and IDT [46] for raindrop; DesnowNet [9], JSTASR [61], DDMSNet [31], and MPRNet [85] for snow.
  • General Image Restoration Methods retrained on the All-Weather dataset: MPRNet* [85], Restormer* [10], and Uformer* [11].
  • UAWR methods: All-in-one [14], TransWeather [15], TKL&MR [17], WeatherDiff [27], AWRCP [19], Tan et al[25], Art [86], T3-DiffWeather [26], and GridFormer [87]
TABLE II: Quantitative Results under Real-World Scenarios. Text Denotes the Best Performance and Underline Denotes the Second Best. The FLOPs are computed on the input image of size $256\times 256$.
Methods SPA [88] RainDS [28] Snow100K-real [9] Params. GFlops
[…]
TransWeather [15] 27.96 0.8677 45.83 21.78 0.690 49.64 60.56 0.388 3.042 37.7 6.13
WeatherDiff128 [27] 29.26 0.9279 46.32 22.40 0.6405 59.00 62.48 0.396 2.970 85.6 65.9$\times$169$\times$25
T3-DiffWeather [26] 31.36 0.9340 45.01 23.12 0.6461 45.01 61.65 0.397 2.910 69.4 59.8$\times$2
T3-DiffWeather (4 steps) [26] 31.65 0.9367 46.10 23.17 0.6388 43.58 61.70 0.397 2.922 69.4 59.8$\times$4
GridFormer [87] 31.18 0.9269 46.23 22.89 0.6377 46.85 61.49 0.395 2.912 30.1 251.35
CyclicPrompt (ours) 31.95 0.9347 46.51 23.10 0.6620 50.32 61.72 0.395 2.893 29.9 230.42
Figure 5: Qualitative comparison of different adverse weather removal methods, i.e., TransWeather [15], GridFormer [87], WeatherDiff et al[27], T3-DiffWeather [26], and our CyclicPrompt under real scenarios. Test images are from RainDS[28], SPA-Data[88] and Snow100K-real[9].
For fair comparisons, we evaluate the results of the methods using their official codes or pre-trained weights once provided. Otherwise, we directly cite the reported results in their papers.
Quantitative Comparison. Table I provides the quantitative results of all methods across three adverse weather types. We can see that the task-specific methods, e.g., MPRNet [85] and IDT [46], produce promising results on the weather type when applying their corresponding customized models or even better than several UAWR methods (e.g., All-in-one [14]). However, as shown in Table I, they fail to be generalized to other out-of-distribution weather types, such as MPRNet on raindrops and IDT on snow, where their PSNR and SSIM decrease significantly. When we retrain the general restoration methods on the same All-Weather dataset according to the UAWR setting, one can observe that the resultant models show better generalizability on three different datasets (MPRNet v.s. MPRNet*). Our CyclicPrompt outperforms existing methods by a significant margin across almost all weather types and achieves the best on average. Compared to the recent most SotA method T3-DiffWeather, our method exceeds it over 0.3dB in average PSNR. This is mainly due to our holistic prompt construction and distinctive cyclic prompt learning design, which is preferable for UAWR.
TABLE III: Quantitative Comparisons on Five Tasks. Text and Denote the Best and Performance and Underline Denotes the Second Best.
Methods
Deraining
Rain100L [89]
Dehazing
SOTS [90]
Denoising
BSD68 [91]
Deblurring
GoPro [92]
Enhancement
LOL [93]
Average
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
HINet[CVPR’21][94] 35.67 0.969 24.74 0.937 31.00 0.881 26.12 0.788 19.47 0.800 27.40 0.875
DGUNet[CVPR’22][95] 36.62 0.971 24.78 0.940 31.10 0.883 27.25 0.837 21.87 0.823 28.32 0.891
MIRNetV2[CVPR’22][96] 33.89 0.954 24.03 0.927 30.97 0.881 26.30 0.799 21.52 0.815 27.34 0.875
SwinIR[ICCV’21][97] 30.78 0.923 21.50 0.891 30.59 0.868 24.52 0.773 17.81 0.723 25.04 0.835
Restormer[CVPR’22][10] 34.81 0.962 24.09 0.927 31.49 0.884 27.22 0.829 20.41 0.806 27.60 0.881
NAFNet[ECCV’22][98] 35.56 0.967 25.23 0.939 31.02 0.883 26.53 0.808 20.49 0.809 27.76 0.881
DL[TPAMI’19][99] 21.96 0.762 20.54 0.826 23.09 0.745 19.86 0.672 19.83 0.712 21.05 0.743
TAPE[ECCV’22][100] 29.67 0.904 22.16 0.861 30.18 0.855 24.47 0.763 18.97 0.621 25.09 0.801
AirNet[CVPR’22][66] 32.98 0.951 21.04 0.884 30.91 0.882 24.35 0.781 18.18 0.735 25.49 0.846
IDR[CVPR’23][101] 35.63 0.965 25.24 0.943 31.60 0.887 27.87 0.846 21.34 0.826 28.34 0.893
InstructIR-5D[ECCV’24][102] 36.84 0.973 27.10 0.956 31.40 0.887 29.40 0.886 23.00 0.836 29.55 0.907
DA-CLIP[ICLR’24][23] - - 31.98 0.971 - - 27.03 0.810 22.09 0.810 - -
PromptIR[NeurIPs’24][82] 36.37 0.972 30.58 0.974 31.31 0.880 - - - - - -
NDR[TIP’24][103] 35.42 0.969 28.64 0.962 31.36 0.887 - - - - - -
DiffUIR[CVPR’24][104] 30.35 0.920 31.08 0.935 - - 27.69 0.845 21.93 0.826 - -
CyclicPrompt (ours) 37.03 0.989 33.02 0.983 31.33 0.941 29.42 0.925 21.79 0.837 30.52 0.935
Qualitative Comparison. In Figure 4, we visualize the restoration results of all compared methods. It can be observed that, under various weather conditions, our method can consistently remove the adverse effects of unfavorable weather from corrupted images and preserve plausible textures. Especially, in the rain+fog degradation, our method effectively recovers obscured scenes with more fine-grained details.
Real Scenario Comparison. We also evaluate the performance of our CyclicPrompt in real-world scenarios under different weather conditions and compare it with more recent methods. We use the SPA [88], RainDS [28], and Snow100K-real [9] datasets as the testing sets. We measure the performance using PSNR, SSIM, and a non-reference metric MUSIQ [105]. For Snow100K-real, since the ground truth images are unavailable, we apply the non-reference metrics MUSIQ, MANIIQA [106], and NIQE [107] for quantitative evaluation. Besides, the model complexity measured by parameters (Params., M) and computational costs (GFLOPs, G) is also reported. As illustrated in Table II, Transweather requires the fewest computations but more parameters than CyclicPrompt and suffers from significant performance reduction. T3-DiffWeather performs efficiently but with over 2×2$\times$2 × parameters and slightly lower scores than CyclicPrompt. To make a more intuitive comparison, we increase the iteration step of T3-DiffWeather to 4, ensuring it has similar computational costs (239.2G) to CyclicPrompt. The results reveal that the model does not yield apparent improvements, which still underperforms CyclicPrompt across most metrics. GridFormer shows relatively heavier model complexity but is not adept in real-world weather, which performs worse. Though WeatherDiff128 is comparable to ours, this method involves about 2.5×2.5$\times$2.5 × more parameters and 1200×1200$\times$1200 × more FLOPs. CyclicPrompt achieves competitive performance with obviously smaller model complexity.
In Fig. 5, we illustrate the visual comparisons on real-world datasets. It is evident that both WeatherDiff128 and T3-DiffWeather are unable to suppress the weather interferences, while CyclicPrompt can produce more visually appealing images. Such quantitative and qualitative performance showcases the superior trade-off between the adaptability, efficiency, and generalizability of CyclicPrompt in practical applications.
Generalizing to More Diverse Degradations. To assess the further effectiveness and generalizability of CyclicPrompt, we apply our method to five restoration tasks: deraining, dehazing, motion deblurring, low-light enhancement, and denoising as in [101]. Table III shows the quantitative performance on Rain100L [89] for deraining, BSD68 [91] for denoising, SOTS [90] for dehazing, GoPro [92] for deblurring, and LOL [93] for low-light enhancement. Here, we compare CyclicPrompt with recent state-of-the-art general restoration methods: HINet [94], DGUNet [95], MIRNetV2 [96], SwinIR [97], Restormer [10], NAFNet [98] and all-in-one restoration methods: DL [99], TAPE [100], AirNet [66], IDR [101], InstructIR-5D [102], DA-CLIP [23], PromptIR [82], NDR [103], and DiffUIR [104]. As illustrated, our CyclicPrompt achieves impressive performance across almost all the datasets in PSNR and SSIM, emphasizing its excellent superiority to diverse degradations.
IV-D Ablation Study
In this subsection, we conduct ablation studies to investigate the effects of all the proposed components in CyclicPrompt. All the models are trained for 300K iterations and evaluated on Outdoor-Rain [33], RainDrop [29], and Snow100K-L [9] datasets. We report their average performance for comparison.
Effect of C2P. As discussed in Sec. III-B, the initial C2P consists of weather-specific knowledge, learnable input-conditional vector, and textual prompt. We denote the variant without any prompt as the baseline model. As shown in Table IV, to verify the effectiveness of the initial C2P, we first induce the weather-specific knowledge in the baseline, which exploits the pre-trained CLIP image encoder to extract visual degradation features, prompting the model to improve. The input-conditional vector can be seen as a dynamic prompt for input content, which can calibrate and adapt the weather-dedicated information to complex weather, thereby providing a favorable gain of 0.21dB. Additionally, in Fig. 6, we conduct visual comparisons between the models with and without the input-conditional vectors, where the results validate that the latter can better remove the weather. Interestingly, the textual prompt corresponding to the weather scene is beneficial for restoration as it can incorporate explicit scene contexts and semantics, even if it stems from low-quality observations. Hence, the model with full C2P achieves the best.
TABLE IV: The Effectiveness of C2P.
Baseline Weather-specific Input-conditional Textual PSNR SSIM
Knowledge Vector Prompt
30.43 0.9176
30.47 0.9180
30.68 0.9195
31.07 0.9229
Figure 6: Visual comparisons between the model without and with the input-conditional vectors, where “IC” denotes the input-conditional vector.
TABLE V: Ablation Study on the Number of Input-conditional Vector.
Vector Number
0 4 8 12
PSNR 30.83 30.87 31.18 30.37
SSIM 0.9217 0.9227 0.9251 0.9222
TABLE VI: The Effectiveness of EPM.
Baseline+C2P Weather-free Prompt Residual Prior PSNR SSIM
31.07 0.9229
31.10 0.9244
31.18 0.9251
Number of Learnable Vectors. The input-conditional vector in C2P is of the size $N\times C$. To further study the impact of the learnable vectors for adverse weather removal, we initialize the vectors with different numbers (i.e., $N$), where “0” denotes the model without the vector. The results are summarized in Table V. We can see that the performance boosts gradually as the number increases. This is due to more vectors yielding better capacity for context understanding and weather-related prompt adaption. However, when we increase $N$ from 8 to 12, the model suffers from performance decreases. This can be attributed to the overlarge vectors introducing unnecessary noisy information that overwhelms other important components. Therefore, we set $N=8$ in our experiments.
Effect of EPM. In this work, the EPM aims at leveraging constrained restoration priors to embed weather-free representations into the C2P, constituting a cyclic C2P to further refine the restoration process. To study its effect, we use the baseline model with initial C2P as the reference. As shown in Table VI, the model replaces the visual prompt in the initial C2P with the weather-free prompt from the generated image in the 1st iteration can discover restoration-related information to improve the decoder features, thereby leading to higher scores.
Figure 7: Visualization of the feature maps before and after RPM.
In CyclicPrompt, the residual prior is incorporated to enrich visual characteristics for better detail recovery along with EPM. We now present the experimental results to verify its impact. First, in Table VI, we observe that it assists well the model and enhances the performance notably. Moreover, in Fig. 7, we visualize the residual prior map and corresponding feature maps before and after the RPM. There are two observations: (1) the prior (Fig. 7(b)) is weather-invariant and allows expressing the contextualized information of scenes. (2) By modulating the decoder feature (Fig. 7(c)) with the prior, the model can effectively emphasize the textures and structures (Fig. 7(d)), improving the representation capacity for reconstruction.
V Conclusion
In this paper, we presented CyclicPrompt, a novel cyclic prompt learning framework for universal adverse weather removal. By integrating weather-specific knowledge, textual contexts, and constrained restoration priors, CyclicPrompt effectively addresses the challenges of handling diverse and complex weather degradations. Based on the proposed composite context prompt (C2P) and erase-and-paste mechanism (EPM), which together form a cyclic “Prompt-Restore-Prompt” pipeline. Extensive experimental results demonstrate the superiority of our CycliPrompt under various weather degradations.